jericakayla:

Sara left a bin outside in her garden to collect rainwater. She notices that 1 over 8 gallon of water fills 2 over 3 of the bin. Write and solve an equation to find the amount of water that will fill the entire bin. Show your work. Explain your answer in words.

Shadow:

\[\frac{ \frac{ 1 }{ 8} gallons } { \frac{ 2 }{ 3 }bin} = \frac{ x gallons }{ 1 bin }\] Cross Multiply \[\frac{ 1 }{ 8 } = \frac{ 2 }{ 3 }x\] Multiply by the reciprocal of 2/3, which is 3/2 \[\frac{ 3 }{ 2 } \times \frac{ 1 }{ 8 } = \frac{ 2 }{ 3 }x \times \frac{ 3 }{ 2 } \] Simplify \[\frac{ 3 }{ 16} = x \] They want us to solve for the amount of gallons that will fill up the bin. 3/16 gallons fills a bin

jericakayla:

thanks!